Determine the [H3O+(aq)] of a solution with a pH of 3.45.

Expressed in scientific notation, the answer is in the form a,b × 10–c.


1
Expert's answer
2022-03-08T05:31:01-0500

-log[H3O+(aq)] = 3.45

log[H3O+(aq)] = -3.45

[H3O+(aq)] = 3.548×10-4